### v1

# Carl Størmer and his Numbers

Abstract.
In many proofs of Fermat’s Two Squares Theorem, the smallest least residue solution $x_{0}$ of the quadratic congruence $x^{2}\equiv-1\bmod p$ plays an essential role; here $p$ is prime and $p\equiv 1\bmod 4$. Such an $x_{0}$ is called a Størmer number, named after the Norwegian mathematician and astronomer Carl Størmer (1874-1957). In this paper, we establish necessary and sufficient conditions for $x_{0}\in\mathbb{N}$ to be a Størmer number of some prime $p\equiv 1\bmod 4$. Størmer’s main interest in his investigations of Størmer numbers stemmed from his study of identities expressing $\pi$ as finite linear combinations of certain values of the Gregory-MacLaurin series for $\arctan(1/x)$. Since less than 600 digits of $\pi$ were known by 1900, approximating $\pi$ was an important topic. One such identity, discovered by Størmer in 1896, was used by Yasumasa Kanada and his team in 2002 to obtain 1.24 trillion digits of $\pi$. We also discuss Størmer’s work on connecting these numbers to Gregory numbers and approximations of $\pi$.
Key words and phrases:
Fermat’s two squares theorem, quadratic congruences, Størmer numbers, natural density
1. Introduction and Motivation
Fermat’s Two Squares Theorem, which G. H. Hardy said is “ranked, very justly, as one of the finest of arithmetic” (see [10]) states
Theorem 1.1.
A prime number $p$ can be written as the sum of squares of two positive integers only when $p\equiv 1\bmod 4.$ Furthermore, in this case, there exists unique positive integers $a$ and $b$ such that
$p=a^{2}+b^{2}.$
From a historical point of view, L. E. Dickson [5] attributes the statement of the above theorem to A. Girard in 1632. In a letter to Mersenne, Fermat’s announced the proof in 1640; E. T. Bell [3, p. 89] and Dickson [7, p. 228] say that Fermat had an irrefutable proof by the method of infinite descent. However, no record of Fermat’s proof has been found. In 1749, Euler gave the first recorded proof of the existence part of Fermat’s theorem. Bell further writes “It was first proved by the great Euler in 1749 after he struggled, off and on, for seven years to find a proof”. The first proof of uniqueness of the positive integers was given by Gauss in 1801.
Most proofs of the existence of the two squares initially involve a study of the quadratic congruence $x^{2}\equiv-1\bmod p$. For example, Hermite [11] and Serret [14] who, independently in 1848, gave proofs of Fermat’s Two Squares Theorem assuming the least residue solution $x_{0}$ of $x^{2}\equiv-1\bmod p$ is known. Then, using the Euclidean Algorithm, they produced algorithms to find the positive integers $a,b$ satisfying Fermat’s Two Squares Theorem. For further information, see the contribution by Brillhart [4] who gave an account of Serret’s and Hermite’s algorithms and improved both.
In 1855, British mathematician Henry John Stephen Smith gave an elegant existence proof of Fermat’s Two Squares Theorem using the Euclidean Algorithm and elementary theory of determinants and continuants; see [5] and [15]. We briefly describe Smith’s method which also emphasizes the importance of the quadratic congruence $x^{2}\equiv-1\bmod p$. For explicit details of Smith’s proof, see [5] and [15].
For $m\in\mathbb{N}$ and positive integers $q_{1},q_{2},\cdots,q_{m}\in\mathbb{N}$, the continuant $q_{1},q_{2},\cdots,q_{m}]$ of length $m$ is defined by $q_{1}]=q_{1}$ and, for $m>1$, it is the determinant
$q_{1},q_{2},\ldots,q_{m}]=\left|\begin{array}[c]{rrrrrr}q_{1}&1&0&\cdots&0&0\\ -1&q_{2}&1&0&\cdots&0\\ 0&-1&q_{3}&1&\cdots&0\\ \vdots&&&&&\vdots\\ 0&0&0&\cdots&q_{m-1}&1\\ 0&0&0&\cdots&-1&q_{m}\end{array}\right|.$
It is not difficult to see that
$q_{1},q_{2},\cdots,q_{n-1},q_{n}]=[q_{n},q_{n-1},\cdots,q_{2},q_{1}];$
for additional properties of continuants, see [5] and [13]. Continuants arise naturally from applying the Euclidean algorithm to the ratio $s/r$ where $s,r\in\mathbb{N}$ and $r<s$. Indeed, if
$s/r$ $=q_{1}+u/r$
$r/u$ $=q_{2}+v/u$
$\vdots$
$x/y$ $=q_{n}+0,$
then
$s=[q_{1},q_{2},\cdots,q_{n}]\quad\text{and}\quad r=[q_{2},q_{3},\cdots,q_{n}].$
Smith proves that if $p\equiv 1\bmod 4$, then there exists an integer $n$ and a unique $x_{0}\in\mathbb{N}$ satisfying $1<x_{0}\leq(p-1)/2$ such that
$p=[q_{1},q_{2},\cdots,q_{n-1},q_{n},q_{n},q_{n-1},\cdots,q_{2},q_{1}]$
and
(1.1) $x_{0}=[q_{2},q_{3}\cdots,q_{n-1},q_{n},q_{n},q_{n-1},\cdots,q_{2},q_{1}].$
Notice that this continuant representation of $p$ is palindromic. The integer $x_{0}$ is the smallest least residue of $x^{2}\equiv-1\bmod{p}$. Furthermore, using elementary properties of determinants, Smith shows
$p$ $=[q_{1},q_{2},\ldots q_{n-1},q_{n}][q_{n},q_{n-1},\ldots,q_{2},q_{1}]$
(1.2) $+[q_{1},q_{2},\ldots,q_{n-1}][q_{n-1},\ldots,q_{2},q_{1}]$
$=[q_{1},q_{2},\ldots,q_{n}]^{2}+[q_{1},q_{2},\ldots,q_{n-1}]^{2}.$
Consequently, determining the least residue solution $x_{0}$ of $x^{2}\equiv-1\bmod{p}$ allows us, through the algorithm in (1.2), to explicitly find positive integers $a,b$ satisfying $p=a^{2}+b^{2}$. To illustrate Smith’s argument, consider $p=13$; in this case $x_{0}=5$ as can be seen from the following calculations using the Euclidean Algorithm,
$\frac{13}{5}=\mathbf{2}+\frac{3}{5},\frac{5}{3}=\mathbf{1}+\frac{2}{3},\frac{3}{2}=\mathbf{1}+\frac{1}{2},\frac{2}{1}=\mathbf{2}+0\text{ so }p=[2,1,1,2]\text{ and }x_{0}=5=[1,1,2].$
It is clear that $x_{0}\equiv-1\bmod 13$. Moreover, employing (1.2), we see that
$13=[2,1,1,2]=[2,1][1,2]+[2][2]=[2,1]+[2]^{2}=3^{3}+2^{2}.$
As we will see in the next section, $x_{0}=5$ is called the Størmer number for the prime $p=13$.
The key point in most proofs of Fermat’s Two Squares Theorem is to first identify the positive integer $x_{0}$ satisfying the two conditions $1<x_{0}\leq(p-1)/2$ and $x_{0}^{2}\equiv-1\bmod p$. In practice, however, when the prime $p\equiv 1\bmod 4$ is a large prime, it is difficult to find $x_{0}$; see Tables 1 and 2 below where it seems that there is considerable randomness in the list of Størmer numbers. In this paper, we obtain necessary and sufficient conditions for testing when $x_{0}\in\mathbb{N}$ is a Stormer number and, in the case that it is, we show how to find the associated unique prime number $p$; see Theorem 2.3 below.
The contents of this paper are as follows. In Section 2, we define Størmer numbers and prove several basic facts about them, including our characterization result (see Theorem 2.3) of Størmer numbers. As far as we can tell, the results in Section 2 are new. In Section 3, we give a heuristic/probabilistic ‘proof’ of the natural density of the Størmer numbers; it was conjectured by Everest and Harman [9] in 2008 that this natural density is $\ln(2)$. Section 4 gives a brief biographical sketch of Carl Størmer; Størmer was a well-known Norwegian number theorist who was also recognized for his ground-breaking work in astronomy and his research on the aurora borealis. Lastly, in Section 5, we discuss Størmer’s work on approximating $\pi$ using the MacLaurin/Gregory series for $\arctan(1/x)$, that is,
$t_{x}:=\arctan\left(\frac{1}{x}\right)=\sum_{k=0}^{\infty}(-1)^{k}\frac{1}{(2k+1)x^{2k+1}}\quad(|x|\geq 1).$
Of course, it is well known that $t_{1}=\pi/4$. Størmer (see [16, 18]) shows that if $x\in\mathbb{N}$ is a non-Størmer number, then $t_{x}$ can be written as a unique finite linear combination of $t_{n}$’s, where each $n$ is a Størmer number (see Theorem 5.1). We illustrate Størmer’s method by considering several examples.
2. Størmer Numbers
Definition 2.1.
Suppose $p=4n+1$ is a prime number for some positive integer $n$. We call a positive integer $x_{0}$ satisfying the two conditions
  1. (i)
    $1<x_{0}\leq\dfrac{p-1}{2}$ (or $3<2x_{0}+1\leq p$)
  2. (ii)
    $x_{0}^{2}\equiv-1\bmod p\vskip 6.0pt plus 2.0pt minus 2.0pt$
a Størmer number for $p$ and we write $S(p)=x_{0}$. If $\mathbb{P}_{4n+1}$ denotes the set of all prime numbers $p\equiv 1\bmod 4$, we call the function $S:\mathbb{P}_{4n+1}\rightarrow\mathbb{N}$, defined by $S(p)=x_{0}$, the Størmer function. We denote the set of Størmer numbers by $\mathbb{S}$ and its complement in $\mathbb{N}$ by $\mathbb{S}^{\text{c}}.$
From the example in the last section, note that $S(13)=5$.
Remark 2.1.
Conway and Guy’s definition of a Størmer number (see [6, p. 245]) is slightly different from ours; they define a Størmer number to be a positive integer $n$ for which the largest prime factor $p$ of $n^{2}+1$ is at least $2n$ while our definition requires the largest prime factor to be at least $2n+1$. Their definition includes $n=1$ as a Størmer number; otherwise, the two definitions agree.
Størmer numbers are listed as item A002314 in Sloane’s website Online Encyclopedia of Integer Sequences.
Table 1 below gives a list of a few ordered pairs $(p,S(p))$ in increasing order of $p$, where the prime $p\equiv 1\bmod 4$ and $S(p)$ is the corresponding Størmer number.
$(5,2)$ $(13,5)$ $(17,4)$ $(29,12)$ $(37,6)$ $(41,9)$ $(53,23)$
$(61,11)$ $(73,27)$ $(89,14)$ $(97,22)$ $(101,10)$ $(109,33)$ $(113,15)$
$(137,37)$ $(149,44)$ $(157,28)$ $(173,80)$ $(181,19)$ $(193,81)$ $(197,14)$
$(229,107)$ $(233,89)$ $(241,64)$ $(257,16)$ $(269,82)$ $(277,60)$ $(281,53)$
$(293,138)$ $(313,25)$ $(317,114)$ $(337,148)$ $(349,136)$ $(353,42)$ $(373,104)$
Table 1
For example, $S(157)=28$ and $S(353)=42$. From this table, it appears that the relationship between $S(p)$ and $p$ is chaotic. As discussed in the introduction, we develop necessary and sufficient conditions for when a given $x_{0}\in\mathbb{N}$ is the Størmer number for some prime $p\in\mathbb{P}_{4n+1}$.
We list the first few Størmer numbers in increasing order:
11 22 44 55 66 99 $10$ $11$ $12$ $14$ $15$ $16$ 1919 2020 $22$
$23$ $24$ $25$ $26$ $27$ $28$ $29$ $33$ $34$ $35$ $36$ $37$ $39$ $40$ $42$
$44$ $45$ $48$ $49$ $51$ $52$ $53$ $54$ $56$ $58$ $59$ $60$ $61$ $62$ $63$
$64$ $65$ $66$ $67$ $69$ $71$ $74$ $77$ $78$ $79$ $80$ $81$ $82$ $84$ $85$
$86$ $87$ $88$ $90$ $92$ $94$ $95$ $96$ $97$ $101$ $102$ $103$ $104$ $106$ $107$
Table 2
Remark 2.2.
For a prime $p,$ it is well known by Euler’s Criterion (see [1, Theorems 9-1 and 9-5]) that the quadratic congruence
(2.1) $x^{2}\equiv-1\bmod p$
has exactly two least residue solutions when $p\equiv 1\bmod 4$ (and no solutions when $p\equiv 3\bmod 4$). When $p\equiv 1\bmod 4,$ one of these solutions lies in the interval (1,(p1)/2](1,(p-1)/2] and the other solution $y_{0}=p-x_{0}$ lies in the interval ((p1)/2,p)((p-1)/2,p). Moreover, it is straightforward to see that $x_{0}=(p-1)/2$ satisfies (2.1) only when $p=5$ (with $x_{0}=2).$ Hence, for primes $p>5$, which are congruent to $1\bmod 4,$ its Størmer number $x_{0}$ satisfies the strict inequality $1<x_{0}<(p-1)/2.$
Our first result is that Størmer’s function is injective.
Theorem 2.1.
The Størmer function $S:\mathbb{P}_{4n+1}\rightarrow\mathbb{N}$, defined in Definition 2.1, is 1-1; that is, if $x_{0}=S(p)$ for some prime number $p,$ then $x_{0}$ cannot be the Størmer number of any other prime $p^{\prime}\in\mathbb{P}_{4n+1}.$
Proof.
Suppose $S(p_{1})=S(p_{2})$ but suppose $p_{1}\neq p_{2}.$ Suppose $p_{1}>p_{2}.$ Since $p_{j}|(x_{0}^{2}+1)$ and we assume $p_{1}\neq p_{2},$ we see that $p_{1}p_{2}|(x_{0}^{2}+1).$ Hence
(2.2) $x_{0}^{2}+1\geq p_{1}p_{2}.$
Furthermore, since $p_{j}$ prime, we see that $p_{2}\leq p_{1}-2.$ Since $x_{0}\leq(p_{2}-1)/2,$ we see that
$x_{0}^{2}$ $\leq(p_{2}-1)^{2}/4$
$=\left(p_{2}^{2}-2p_{2}+1\right)/4$
$\leq\left(p_{2}(p_{1}-2)-2p_{2}+1\right)/4$
$=(p_{1}p_{2}-4p_{2}+1)/4$
and hence
(2.3) $4x_{0}^{2}-1\leq p_{1}p_{2}-4p_{2}<p_{1}p_{2}.$
However, $x_{0}^{2}+1<4x_{0}^{2}-1$ so ((2.3)) implies that
(2.4) $x_{0}^{2}+1<p_{1}p_{2}.$
However ((2.4)) contradicts ((2.2)) unless $p_{1}=p_{2}.$ Consequently, the Størmer map is 11-11. ∎
It follows from Theorem 2.1 that there are infinitely many Størmer numbers; i.e. $|\mathbb{S}|=\aleph_{0}$, where $|A|$ denotes the cardinality of a set $A$.
For $x_{0}\in\mathbb{N}$, suppose the prime factorization of $x_{0}^{2}+1$ is
(2.5) $x_{0}^{2}+1=2^{r}p_{1}^{r_{1}}p_{2}^{r_{2}}\cdots p_{m}^{r_{m}},$
where $r$ is a non-negative integer, each $r_{j}\in\mathbb{N},$ and each $p_{j}$ is an odd prime number with $p_{1}<p_{2}<\ldots<p_{m-1}<p_{m}.$
Theorem 2.2.
Suppose $n=x_{0}^{2}+1$ has the prime factorization given in (2.5). Then, for each $1\leq j\leq m,$ $p_{j}\equiv 1\bmod 4.$
Proof.
Suppose, to the contrary, that $p_{j}\equiv 3\bmod 4$ for some $1\leq j\leq m.$ Since $p_{j}|(x_{0}^{2}+1),$ we see that $x_{0}^{2}\equiv-1\bmod p_{j}.$ Furthermore, $\gcd(x_{0},p_{j})=1$ (otherwise $p_{j}|x_{0}$ and $p_{j}|(x_{0}^{2}+1)$ so $p|1$ which is not possible), we see by Fermat’s Little Theorem that
(2.6) $x_{0}^{p_{j}-1}\equiv 1\bmod p_{j}.$
Then, since $(p_{j}-1)/2$ is odd, we see that
$x_{0}^{p_{j}-1}=(x_{0}^{2})^{(p_{j}-1)/2}=(-1)^{(p_{j}-1)/2}\equiv-1\bmod p_{j},$
which contradicts (2.6). ∎
We are now in position to characterize which positive integers $x_{0}$ can be the Størmer number of some prime $p\equiv 1\bmod 4$.
Theorem 2.3.
Suppose $n=x_{0}^{2}+1$ has the prime factorization given in (2.5) where we assume the primes satisfy $p_{1}<p_{2}<\ldots<p_{m}$. Then there exists a prime $p$ for which $x_{0}$ is the Størmer number if and only if $2x_{0}+1\leq p_{m}$. Moreover, if this condition is met, then $p=p_{m}$ and $S(p_{m})=x_{0}.$ Otherwise, $x_{0}$ is not the Størmer number for any prime $p\equiv 1\bmod 4.$
Proof.
Suppose $2x_{0}+1\leq p_{m};$ that is,
$x_{0}\leq\frac{p_{m}-1}{2}.$
Since $p_{m}|(x_{0}^{2}+1),$ we see that $x_{0}$ is the Størmer number for $p=p_{m};$ i.e. $S(p_{m})=x_{0}.\vskip 6.0pt plus 2.0pt minus 2.0pt$
Conversely, suppose $x_{0}=S(p)$ for some prime $p\equiv 1\bmod 4.$ It follows from the prime factorization of $x_{0}^{2}+1$ that $p=p_{j}$ for some $j\in\{1,2,\ldots m\}$. By definition, $2x_{0}+1\leq p_{j}$ and $x_{0}^{2}\equiv-1\bmod p_{j}.$ Suppose, for the sake of contradiction, that $j<m$. Then since $p_{m}>p_{j}$ $\geq 2x_{0}+1$ and $x_{0}^{2}\equiv-1\bmod p,$ we see that $x_{0}$ is also the Størmer number for $p_{m}.$ But this contradicts the fact that $S$ is 1-1. This forces $p=p_{m}.$
If $2x_{0}+1>p_{m},$ then $x_{0}>(p_{m}-1)/2$ so, by definition $x_{0}$ is not the Størmer number for $p_{m}$ nor, as the above argument shows, can it be the Størmer number for any other prime $p\equiv 1\bmod 4.$
Examples
  1. (1)
    Let $x_{0}=3.$ Since $x_{0}^{2}+1=3^{2}+1=2\cdot 5$ and $2x_{0}+1=7>5,$ $x_{0}=3$ is not the Størmer number for any prime $p\equiv 1\bmod 4.\vskip 3.0pt plus 1.0pt minus 1.0pt$
  2. (2)
    Let $x_{0}=15.$ Since $x_{0}^{2}+1=226=2\cdot 113$ and $2x_{0}+1=31\leq 113,$ it is the case that $S(113)=15$ (see Table 1).
  3. (3)
    Let $x_{0}=279.$ In this case, $x_{0}^{2}+1=2\cdot 38921$ and since $2x_{0}+1=559\leq 38921,$ we see that $x_{0}=279$ is the Størmer number for $p=38921$.
  4. (4)
    Suppose $p$ is a prime of the form $4n^{2}+1$ (for example, $5,17,$ or $37).$ Then $x_{0}=2n$ is the Størmer number for $p=4n^{2}+1$ since $2x_{0}+1=4n+1\leq 4n^{2}+1.$
3. Comments on the Natural Density of Størmer Numbers
The following table seems to suggest that, as $n\rightarrow\infty,$ the number $n$ of positive integers which are Størmer numbers nears $70\%.$
First $n$ positive integers # of Størmer numbers
$100$ $86$
$1000$ $719$
$10000$ $7101$
$100000$ $70780.$
$1000000$ $704536$
Table 3
In fact, Everest and Harman [9, Conjecture 1.5] conjecture
(3.1) $\lim_{n\to\infty}\dfrac{|\{S(p)\mid p\leq n\}|}{n}=\ln 2.$
We now give a heuristic/probabilistic argument of (3.1). We note that a rigorous proof still remains elusive. Let $x_{0}\in\mathbb{N}$. If $x_{0}=S(p)$ for some prime $p\equiv 1\bmod 4$, then $p\geq 2x_{0}+1$. Further, since $x_{0}^{2}\equiv 1\bmod p$ and $p\mid(x_{0}^{2}+1)$, we see that $p\leq x_{0}^{2}+1$. Thus, for any prime $p\equiv 1\bmod 4$, $S(p)$ can only be the Størmer number for $p$ if we have
$2x_{0}+1\leq p\leq x_{0}^{2}+1.$
Assuming each of the integers from $1$ to (p1)/2(p-1)/2 is equally likely to be the Størmer number for $p$, we can informally think of the “probability” that $x_{0}$ is a Størmer number for the prime $p$ as the reciprocal of (p1)/2(p-1)/2, that is, 2/(p1)2/(p-1). Then since $S(p)$ is injective, so that $x_{0}$ is a Størmer number for at most one prime, we can sum these all up to see that the “probability” that $x_{0}$ is a Størmer number for any prime $p$ is
$\sum_{\begin{subarray}{c}2x_{0}+1\leq p\leq x_{0}^{2}+1\\ p\equiv 1\bmod 4\end{subarray}}\frac{2}{p-1}.$
This sum behaves like
$\sum_{\begin{subarray}{c}2x_{0}\leq p\leq x_{0}^{2}\\ p\equiv 1\bmod 4\end{subarray}}\frac{2}{p}$
since the effect of altering $p$ by 1 goes to zero in the limit. Since in the limit, there are the same number of primes of the form $p\equiv 1\bmod 4$ and $p\equiv 3\bmod 4$ (due to the Prime Number Theorem for arithmetic progressions), we write this as half the sum over all primes – not just the primes of the form $p\equiv 1\bmod 4$ – to obtain
$\sum_{2x_{0}\leq p\leq x_{0}^{2}}\frac{1}{p}.$
Then using Merten’s estimate (see [2, Lemma 4.10]) we obtain
$\sum_{p\leq x_{0}}\frac{1}{p}\sim\ln\ln x_{0},$
we write this asymptotic as
$\sum_{p\leq x_{0}^{2}}\frac{1}{p}-\sum_{p<2x_{0}}\frac{1}{p}\sim\ln(\ln x_{0}^{2})-\ln(\ln 2x_{0})\sim\ln\left(\frac{\ln x_{0}^{2}}{\ln 2x_{0}}\right)\sim\ln\left(\frac{2\ln x_{0}}{\ln x_{0}}\right)\sim\ln 2.$
The gap in rigor in this argument is, of course, the introduction of probability and the assumption that $x_{0}$ has a 2/(p1)2/(p-1) chance of being the Størmer number for any given prime $p$.
4. A Biographical Sketch of Carl Størmer
Figure 1. Carl Størmer
Carl Størmer (1874-1957) was a preeminent Norwegian mathematician, botanist, and astrophysicist. For his lifetime contributions to mathematics and astronomy, particularly his study of the aurora borealis and the motion of charged particles in magnetic fields, he was bestowed several honors. A crater on the far side of the Moon is named after him. He was elected to several Scandinavian academies, the Royal Society of London and the Paris Academy of Sciences. He was given honorary degrees by the universities of Oxford, Copenhagen and the Sorbonne. The Paris Academy of Sciences awarded him their Janssen Medal in 1922. He was invited to give a one-hour lecture, on his scientific research on the aurora borealis, at the International Congress of Mathematicians in Toronto in 1924 and, in 1936, he was president of the International Congress of Mathematicians which was held in Oslo.
Størmer entered the University of Christiana (Oslo’s previous name) in 1892 and obtained his candidates’ degree (similar to a Ph.D. degree) in 1897. By this time, Størmer had written several mathematical articles, primarily in number theory, as well as a number of short notes in botany. From 1898-1900, Størmer studied at the Sorbonne under mathematical giants Émile Picard, Henri Poincaré, Camille Jordan, Gaston Darboux, and Édouard Goursat. His output of mathematical papers continued with twelve papers on series, number theory, and the theory of functions between 1896 and 1902. He worked with Sylow and Holst to produce two volumes in 1902 to celebrate the centenary of Niels Henrik Abel’s (1802-1829) birth.
Between 1896-1899, Størmer published two papers on connecting Størmer numbers to Gregory numbers, the subject of this paper. In 1897, he proved that if $P$ is a finite set of prime numbers, then there are only finitely many consecutive integers having only the numbers from $P$ as their prime factors. Moreover, Størmer develops an algorithm for finding all such pairs of consecutive integers. To prove his result, he reduced the problem to solving a finite number of Pell’s equations. This paper [17] was highly praised by the number theorist Louis Mordell.
In 1903, Størmer was appointed professor of pure mathematics at the University of Oslo, a position he held until 1946. In 1934, Størmer published a paper in the Norwegian Mathematical Society entitled ‘Ramanujan - A Remarkable Genius’. At the insistence of fellow Norwegian Atle Selberg, a leading authority in number theory, this paper was reprinted in the same journal in 1989. Selberg was a teenager in 1934 and remarked that Størmer’s paper on Ramanujan contributed to arousing his interest in mathematics.
During his time at the University of Oslo, Størmer’s interests were also drawn to another scientific area: the study of the aurora borealis. A colleague of Størmer, Kristian Berkeland, had put forward a theory in 1896 that auroras were caused by electrons emitted by the sun which interacted with the earth’s magnetic field. Poincaré had, in the same year, solved the differential equations resulting from the motion of a charged particle in the field of a single pole. This was not the situation for an aurora since the magnetic field of the earth is a dipole. Størmer specifically attacked this problem. His papers on this subject constitute a major part of his life’s work. Størmer wrote two books, From the Depths of Space to the Heart of the Atom, which was translated into five languages, and The Polar Aurora which contains both his experimental work on aurorae and his mathematical attempts to model them. The 2013 book Carl Størmer: Aurora Pioneer, by Egeland and Burke [8] is a comprehensive account of Størmer’s contributions to the study of auroras.
We end this sketch with a rather humorous note about Carl Størmer. He was an avid photographer which he put to use in his study of auroras. He purchased a miniature spy camera that he carefully concealed in his jacket. He would walk the streets of Oslo and take photographs of people. He published two works of his ‘snapshots of famous people’ in 1942 and 1943. Long afterwards, nearing the age of 70, these photographs formed the subject of a major exhibition in Oslo. Many of these photographs exist on the internet and they capture a fascinating look at life in Oslo during the early part of the 20th century.
5. Application of Størmer Numbers: Gregory Numbers and Approximations of $\pi$
As of this writing, there are more than $200$ trillion digits of $\pi$ known. Indeed, using A. Yee’s y-cruncher (see [21]) computer program, the StorageReview Lab Team computed the first 202,112,290,000,000 digits of $\pi$ in 2024; see [19]. However, in the early 1900’s, there were less than $600$ digits of $\pi$ known. Of course, manual calculations were the only option then to approximate $\pi$ long before the modern computer age. Archimedes, circa 250 BC, showed that
$\frac{223}{71}<\pi<\frac{22}{7}$
Zu’s ratio, discovered by the Chinese astronomer Zu Chongzhi in the 5th century, gives
$\pi\approx\frac{355}{113}\approx 3.14159292035,$
an approximation within $0.000009\%$ of the value of $\pi$; in fact, Zu’s approximation of $\pi$ is the best rational approximation to $\pi$ with a denominator of four or fewer digits.
In the $18^{\text{th}}$ century, due to the work of John Machin and Leonhard Euler, techniques were developed to approximate $\pi$ using Gregory numbers $t_{x}$, defined by
$t_{x}:=\arctan\left(\frac{1}{x}\right)=\sum_{k=0}^{\infty}(-1)^{k}\frac{1}{(2k+1)x^{2k+1}}\qquad(|x|\geq 1).$
Of course, it is well-known that that $t_{1}=\pi/4$; this particular series is known as the Mādhava-Gregory-Leibniz series. Convergence is slow; indeed, roughly speaking, one must take 10 times more terms to add one extra decimal place of accuracy in the computation of $\pi$. Euler established
(5.1) $t_{1}=5t_{7}+2t_{79/3},$
while Machin proved that
(5.2) $t_{1}=4t_{5}-t_{239}$
which, of course, is equivalent to the well-known trigonometric identity
(5.3) $\pi=16\arctan\left(\frac{1}{5}\right)-4\arctan\left(\frac{1}{239}\right).$
Convergence is faster with Machin’s formula; indeed, the first 100 terms of the series on the right-hand side of Machin’s formula (5.2) produces the first 140 digits of $\pi$.
In the late $19^{\text{th}}$ century, Carl Størmer noticed an important connection between Gaussian primes and identities involving Gregory numbers. It is the purpose of this section to discuss his work; the mathematics behind the technique he develops is fascinating and deserves attention certainly not because of its application to approximating $\pi$ in today’s technological world but for its sheer mathematical elegance and beauty. For further information on Størmer’s work on Gregory series and approximating $\pi$, see Størmer’s work in [16, 18] as well as the book by Conway and Guy [6].
Among his discoveries, in 1896, Størmer proved the identity
(5.4) $t_{1}=44t_{57}+7t_{239}-12t_{682}+24t_{12943}.$
In 2002, the Japanese computer scientist Yasumasa Kanada and his team used this identity to help obtain the first $1.2411$ trillion digits of $\pi$ which, at the time, was the world record for most digits of $\pi$. We next discuss Størmer’s theorem connecting Gregory series with Størmer numbers; for a reference, see [6] and [16].
Theorem 5.1 (Størmer).
If $x\in\mathbb{N}$ is not a Stormer number, then $t_{x}$ can be uniquely expressed as a finite linear combination of $t_{n}$’s where each $n$ is a Størmer number and each $n<x$.
We note that in Machin’s formula (5.2), $239$ is not a Størmer number but both 1 and 5 are while, in Størmer’s formula (5.4), none of the integers 57, 239, 682, 12943 are Størmer numbers.
Recall that a complex number $z=a+bi$, where $a,b\neq 0$, is a Gaussian prime if $a^{2}+b^{2}$ is prime. For example, $i\pm 1$ and $3i\pm 2$ are Gaussian primes but $3\pm i$ is not. To illustrate Størmer’s work, we discuss an important relationship between Gregory numbers $t_{x}$, when $x\in\mathbb{N}$, and Gaussian primes. If $z=a+bi$, where $a>0$ and $\theta=\arg(z)$, where $\arg(z)\in(-\pi/2,\pi/2)$, then $\tan\theta=b/a$ showing that $\theta=\arctan(b/a)$. In particular, for $n\in\mathbb{N}$, we see that
(5.5) $\arg(n+i)=\arctan\left(\frac{1}{n}\right)=t_{n}.$
Moreover, since $\arg(z_{1}z_{2})=\arg(z_{1})+\arg(z_{2})$, we see, for $c>0$ and $k\in\mathbb{N}$, that
(5.6) $\arg(cz)$ $=\arg(z)$
(5.7) $\arg(z^{k})$ $=k\arg(z).$
Now suppose that $t_{n}=t_{n_{1}}+t_{n_{2}}$, where $n,n_{1},n_{2}\in\mathbb{N}$. Then, from (5.5),
(5.8) $\arg((n_{1}+i)(n_{2}+i))=t_{n_{1}}+t_{n_{2}}=t_{n}.$
With an abuse of notation, but for better readability, we shall write (5.8) as
(5.9) $n+i=(n_{1}+i)(n_{2}+i).$
This is the same notation adopted by Conway and Guy [6].
In order to use Theorem 5.1, we must make repeated use of the identities in (5.5), (5.6), (5.7), and (5.8). We describe how we do this in the next few paragraphs.
Suppose $n\in\mathbb{S}^{\text{c}}$; that is, $n$ is not a Størmer number. To determine the representation of $t_{n}$ guaranteed by Størmer’s Theorem, we first consider the Gaussian prime factorization of $(n+i)$. Within this factorization, there are four simple forms that factors may assume:
  1. (I)
    $m\pm i$ with $m\in\mathbb{S}$;
  2. (II)
    $a+bi$ with $a,b\in\mathbb{Z}$ and $\gcd(a,b)=1$ (which we may assume from (5.6));
  3. (III)
    $l\pm i$ with $l\in\mathbb{S}^{\mathsf{c}}$;
  4. (IV)
    $\pm i$.
Because equality depends only on the argument, all positive scalars in the Gaussian prime factorization may be cast aside as they do not affect the argument of the terms.
These rules aim to maintain the $(n+i)$ term and convert all other terms in the prime factorizaton to factors of the form (I). To aid in this process there are a few key identities that are useful in our calculations:
  1. (a)
    $(1-i)^{4}=(1+i)^{4}=-1$
  2. (b)
    $(1+i)^{7}=(1-i)$
  3. (c)
    $(1-i)^{7}=(1+i)$
  4. (d)
    $i=(1+i)^{2}$
  5. (e)
    $i=(1-i)^{2}$
  6. (f)
    i(a+bi)=(b+ai)i(a+bi)=(-b+ai);
again, equality in these ‘identities’ is in the sense of (5.8), namely up to positive integer multiples.
To simplify terms of form (II), we ‘flatten’ a complex number $a+bi$ by replacing it by a complex number of the form $x\pm i$. To do this, we find $c,d,e\in\mathbb{Z}$ such that $(a+bi)(c+di)=(e\pm i)$. Expanding and separating the real and imaginary components, this equation yields the system of equations:
$ac-bd$ $=e$
$ad+bc$ $=\pm 1.$
In particular, we must find a solution $(c,d)$ to the linear Diophantine equation $ad+bc=1$. Ideally, $d=\pm 1$, so $(c+di)$ is of the form (I). However, in many cases as we illustrate below, this process of flattening might need to be repeated several times.
Example 5.1.
The complex number 185i18-5i is of the form (II). To reduce it to a complex number of form (I), we solve the Diophantine equation 18d5c=118d-5c=1 to obtain $(c,d)=(7,2)$ which corresponds to $7+2i$. Since this latter number is not of the form (I), we solve $7d+2c=1$ to obtain (c,d)=(4,1)=4i(c,d)=(4,-1)=4-i which is of the form (I).
By calculating the Størmer expansions of Gregory numbers in increasing order, terms of form (III) may be substituted with their expanded forms. Finally, terms of form (IV) may be dealt with via our above key identities.
We note that others have studied Størmer numbers in their connection to the arctangent or cotangent functions. Since $\arctan(1/x)=\operatorname{arccot}(x)$, Lehmer [12] studied identities of the form
$\operatorname{arccot}(a/b)=\operatorname{arccot}(n_{0})-\operatorname{arccot}(n_{1})+\operatorname{arccot}(n_{2})$-$\cdots$
where the integers $n_{j}$ are obtained by solving the recurrences
$a_{j}=n_{j}b_{j}+b_{j+1}\quad(0\leq b_{j+1}<b_{j}),\quad a_{j+1}=a_{j}n_{j}+b_{j}$
with initial conditions $a_{0}=a,b_{0}=b$. Todd (see [20] and the references cited therein) uses the notation $(n)=\arctan(n)$ instead of $t_{n}=\arctan(1/n)$ so, in our language, $(n)=t_{1/n}$. Todd defines, for $n\in\mathbb{N}$, ($n$) to be irreducible if it cannot be written as a finite sum of the form
$\sum\alpha_{j}(n_{j}),$
where each $\alpha_{j}\in\mathbb{Z}$; otherwise, $(n)$ is reducible. He proves the following theorem (see [20, Theorem B]) which is a slightly weaker version of Theorem 5.1.
Theorem 5.2.
For $\in\mathbb{N}, nn$ is a Stormer number if and only if $(n)$ is irreducible.
In [20, Theorem A], Todd also proves the following characterization theorem:
Theorem 5.3.
For $\in\mathbb{N}, nn$ is a Stormer number if and only if all prime factors of $1+n^{2}$ should occur among the prime factors of $1+m^{2}$ for $m=1,2,\cdots n-1$.
In the examples that follow, we use our notation to find identities involving Gregory numbers $t_{n}$.
Example 5.2.
The Gaussian prime factorization of $3+i$ is given by
3+i=i(1+i)(1+2i)=(1+i)(2i)3+i=-i(1+i)(1+2i)=(1+i)(2-i)
so $t_{3}=t_{1}-t_{2}$. We remark that, using Todd’s notation, this formula is equivalent to (3)=3(1)(2)(3)=3(1)-(2).
Example 5.3.
Vega’s identity is
(5.10) $t_{1}=2t_{3}+t_{7}.$
The Gaussian prime factorizations of $3+i$ and $7+i$ are
(5.11) $3+i$ $=-i(1+i)(1+2i)=(1+i)(2-i)$
(5.12) $7+i$ $=-i(1+i)(2+i)^{2}=(1-i)(2+i)^{2}.$
From (5.11), we see that $t_{3}=t_{1}-t_{2}$ while (5.12) yields $t_{7}=-t_{1}+2t_{2}$. Combining these two identities gives us (5.10).
Example 5.4.
The number $n=21$ is not a Størmer number by Theorem 2.3 since $21^{2}+1=2\cdot 13\cdot 17$ and $2n+1=43>17$. The Gaussian Prime factorization of $21+i$ is,
(5.13) 21+i=(i)(1+i)(3+2i)(4+i).21+i=(-i)(1+i)(3+2i)(4+i).
Both $1$ and $4$ are Størmer numbers so we only need to convert i-i and $3+2i$. A solution of the equation 3d+2c=13d+2c=-1 is (c,d)=(1,1)=1i(c,d)=(1,-1)=1-i. Multiplying both sides of (5.13) by 1i1-i, we find that
(1i)(21+i)=(i)(1+i)(5i)(4+i).(1-i)(21+i)=(-i)(1+i)(5-i)(4+i).
Next we use key identity (e) above, namely $(-i)=(1-i)^{2}$, to obtain
$(1-i)(21+i)=(1-i)^{2}(1+i)(5-i)(4+i).$
To isolate $21+i$, we multiply both sides by $1+i$ and obtain
$21+i=(1-i)^{2}(1+i)^{2}(5-i)(4+i)=(5-i)(4+i)$
and this yields the identity $t_{21}=t_{4}-t_{5}$.
Example 5.5.
Since $70^{2}+1=13^{2}\cdot 29$, Theorem 2.3 says that $n=70$ is not a Størmer number. The Gaussian prime factorization of $70+i$ is
(5.14) $70+i=-(2+3i)^{2}(2+5i).$
We first eliminate the $2+5i$ term. By inspection, a solution to $2d+5c=1$ is (c,d)=(1,2)(c,d)=(1,-2) so multiplying (5.14) by 12i1-2i yields
(5.15) $(1-2i)(70+i)=-(2+3i)^{2}(12+i).$
Solving $2d+3c=1$ produces the solution (c,d)=(1,1)(c,d)=(1,-1); multiplying (5.15) by $(1-i)^{2}$ yields
$(1-i)^{2}(1-2i)(70+i)=-(5+i)^{2}(12+i).$
Finally, multiplying both sides by $i=(1-i)^{2}$ (see key identity (d)), we obtain
$(1-i)^{2}(2+i)(70+i)=-(1-i)^{2}(5+i)^{2}(12+i),$
which yields the identity $t_{2}+t_{70}=2t_{5}+t_{12}$ or $t_{70}=-t_{2}+2t_{5}+t_{12}$.
Example 5.6.
We use Størmer’s technique to establish Euler’s identity in (5.1). We begin by flattening $79+3i$; a calculation shows that
(5.16) (79+3i)(26+i)=2057+i.(79+3i)(-26+i)=-2057+i.
The Gaussian prime factorization of 2057+i-2057+i is
$-2057+i$ $=i(1+i)(1+2i)^{5}(1+26i)$
(5.17) $=(1+i)(1+2i)^{5}(-26+i).$
Substituting (5.17) into (5.16) yields
(5.18) $(79+3i)(-26+i)=(1+i)(1+2i)^{5}(-26+i),$
or
(5.19) $79+3i=(1+i)(1+2i)^{5}.$
We now flatten $1+2i$ by solving $d+2c=1$ to obtain (c,d)=(1,3)(c,d)=(-1,3). Calculations show $(-1+3i)^{5}=-(316+12i)=-4(79+3i)$ and $(1+2i)^{5}(-1+3i)^{5}=(-7+i)^{5}$. Hence, multiplying both sides of (5.19) by $(-1+3i)^{5}$ yields
$(79+3i)^{2}=(1+i)(-7+i)^{5}=-(1+i)(7-i)^{5},$
or $2t_{79/3}=t_{1}-5t_{7}$.
Example 5.7.
In our last example, we verify Machin’s identity in (5.2). The Gaussian factorization of $239+i$ is given by
$239+i=i(1+i)(2+3i)^{4}$
Since $i=(1+i)^{2}$, we see that
(5.20) $239+i=(1+i)^{3}(2+3i)^{4}.$
To eliminate $2+3i$, we solve $2d+3c=1$ to obtain (c,d)=(1,1)(c,d)=(1,-1). Since (1i)(2+3i)=5+i(1-i)(2+3i)=5+i, we multiply both sides of (5.20) by $(1-i)^{4}$ to obtain
(5.21) $(239+i)(1-i)^{4}=(1+i)^{3}(5+i)^{4}.$
Lastly, multiply both sides of (5.21) by $(1+i)^{4}$ and employing key identity (c) yields
$239+i=(1+i)^{7}(5+i)^{4}=(1-i)(5+i)^{4},$
and so $t_{239}=-t_{1}+4t_{5}$.
Note: Graeme Reinhart is an undergraduate student majoring in mathematics at Baylor University.
Author Contributions: Matthew Kroesche, Lance Littlejohn, and Graeme Reinhart equally contributed to the results in this manuscript and each of these authors helped in the final editing process. All authors agree to be accountable for all aspects of the work.
Disclosure of interest: The authors have no competing interests to declare.
Funding: No funding was received.